2) On ONE sketch, show the indifference curves through the points (1,1) and (2, 4) for the following utility functions (a) ua (x, y)-y (b) ub(x, y) -zy (d) uy) 3,,3
Answer #1

Answer:

A) U=x*y

For (1,1), U=1*1=1

So equation for indifference curve x*y=1 for (1,1)

For (2,4), U=2*4=8

So equation for indifference curve x*y=8 for (2,4)

So Graph is

Indifference Curve 9.00 8.00 7.00 6.00 5.00 4.00 3.00 U-8 1.00 U=1 0.00 1 2 345 6 78910

B)

U=xy^2

For (1,1), U=1*1^2=1

So equation for indifference curve x*y^2=1 for (1,1)

For (2,4), U=2*4^2=32

So equation for indifference curve x*y^2=32 for (2,4)

So Graph is

c)

U=x^3y

For (1,1), U=1^3*1=1

So equation for indifference curve x^3*y=1 for (1,1)

For (2,4), U=2^3*4=32

So equation for indifference curve x^3*y=32 for (2,4)

So Graph is

D)

U=x^3y^3

For (1,1), U=1^3*1^3=1

So equation for indifference curve x^3*y^3=1 for (1,1)

For (2,4), U=2^3*4^3=512

So equation for indifference curve x^3*y^3=512 for (2,4)

So Graph is

ndifference Curve 9.00 8.00 7.00 6.00 5.00 4.00 3.00 U#512 U=1 1.00 0.00 1 2 345 6 78910